A ball rolling down an incline travels s feet in t seconds, where s= 5t^2. Derive a formula for the velocity of the ball at time t= t0. How fast is it going (a) after 2 seconds (b) after it has rolled 80 feet?

1
Expert's answer
2022-06-02T16:56:10-0400
"v(t)=s'(t)=(5t^2)'=10t"

"v(t_0)=10t_0"

(a) "t_1=2s"


"v(t_1)=10(2)=20(ft\/s)"

(b)


"s(t_2)=5t_2^2=80, t_2\ge0"

"t_2=\sqrt{16}=4(s)"

"v(t_2)=10(4)=40(ft\/s)"